OpenStudy (anonymous):

Can someone show me how to use the Zero Product Property to find the root of: x^2+6x ?

OpenStudy (unklerhaukus):

do you mean x^2+6x = 0 ?

OpenStudy (anonymous):

Yeah.

OpenStudy (unklerhaukus):

factor out an x from those terms x^2+6x = 0 x( .... ) =0

OpenStudy (unklerhaukus):

can you do this?

OpenStudy (anonymous):

I'm a little confused...

OpenStudy (unklerhaukus):

you know how x^2 is the same as x*x right?

OpenStudy (anonymous):

Yes.

OpenStudy (unklerhaukus):

so you have x^2+6x = 0 x*x+6*x = 0 on the left hand side both terms have a factor of x factor it out like this (ab+cb) = b(a+c)

OpenStudy (unklerhaukus):

x*x+6*x = x( .... + ...)

OpenStudy (anonymous):

I think I'm just confused because there is no value in the middle. Before this question I answered: \[3x ^{2}-7x+4=(x-1)(3x-4)\] I used a generic rectangle to factor it out, and I'm only comfortable doing these with a value like the 4.

OpenStudy (unklerhaukus):

that is factoring a quadratic, this question in not a quadratic, its easier

OpenStudy (anonymous):

Ohhh! Okay, but I'm still confused, could you dumb it up a little bit more?

OpenStudy (unklerhaukus):

you know how to expand brackets like this right? 2(y+3) = 2y+6 the factoring step is the reverse 2y+6 = 2(y+3)

OpenStudy (unklerhaukus):

xx+6x = x( ...+ ...)

OpenStudy (unklerhaukus):

what are the dots?

OpenStudy (anonymous):

x+6?

OpenStudy (unklerhaukus):

yes x^2+6x = 0 x(x+6) =0

OpenStudy (unklerhaukus):

now you have two terms, and their product equals zero for this to be true, one of the factors either x or (x+6) must equal zero

OpenStudy (unklerhaukus):

so x=0 or x+6=0 what values for x satisfy this ?

OpenStudy (anonymous):

Well, 0 for the first one, and -6 for the second?

OpenStudy (unklerhaukus):

Correct!

OpenStudy (unklerhaukus):

so those are the two roots.

OpenStudy (unklerhaukus):

if you were to graph the equation you would get something like this 

OpenStudy (anonymous):

Thank you so much! @UnkleRhaukus